# MoEMeta: Mixture-of-Experts Meta Learning for Few-Shot Relational Learning

### v2

@@ 4.2 MoE-based Meta-Knowledge Learning
$\mathbf{r}_{j}$ $=\frac{1}{N}\sum_{i=1}^{M}g_{i,j}\text{MLP}_{i}(\mathbf{h}_{j},\mathbf{t}_{j};\bm{\theta}_{i}),$ (5)
$g_{i,j}$ $=\begin{cases}s_{i,j},&\text{if }s_{i,j}\in\text{Top}_{j}\left(\{s_{i,j}\mid 1\leq i\leq M\},N\right),\\ 0,&\text{otherwise},\end{cases}$ (6)
$s_{i,j}$ $=\text{softmax}(\text{Gate}(\mathbf{h}_{j},\mathbf{t}_{j};\bm{\theta}_{g})),$ (7)
where $M$ denotes the total number of experts, each implemented as a two-layer MLP, $\text{MLP}_{i}(\cdot;\bm{\theta}_{i})$. For each support triplet $j$, a gate network, $\text{Gate}(\cdot;\bm{\theta}_{g})$, also implemented as a two-layer MLP, computes a relevance score $s_{i,j}$ for each expert $i$. These scores are normalized via a softmax function over all $M$ experts. The function $\text{Top}_{j}(\cdot,N)$ selects only the top $N$ highest scores among all $M$ experts for each triplet $j$, setting the remaining scores to zero. This sparsity mechanism enables the model to selectively activate only a subset of experts for each support triplet, thereby facilitating the learning of more specialized relational prototypes globally shared across tasks.
@@ 4.3 Task-Tailored Local Adaptation
Specifically, for each task, MoEMeta maintains a task-specific set of local parameters, denoted as $\bm{\eta}$, namely three randomly initialized projection vectors, $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$, which project the embeddings $\mathbf{h}_{j}$, $\mathbf{R}_{\mathcal{T}_{r}}$, and $\mathbf{t}_{j}$ into a task-specific subspace. Inspired by TransD (Ji et al., 2015), which addresses the limitations of fixed mapping strategies using per-entity and per-relation projections, our approach leverages task-specific projections to avoid overparametrization under few-shot settings. As all triplets within a task share the same relation, we adopt a unified projection vector $\mathbf{p}_{h}$ for all head entities, and $\mathbf{p}_{t}$ for all tail entities within the task. This provides a lightweight solution for enabling efficient task-specific adaptation with minimal parameter overhead.
@@
$\mathbf{h}_{j}^{\prime}$ $=\mathbf{h}_{j}+(\mathbf{p}_{h}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}},$ (9)
@@
$\mathbf{t}_{j}^{\prime}$ $=\mathbf{t}_{j}+(\mathbf{p}_{t}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}}.$ (11)
Above, the dot product $\mathbf{p}^{\top}\mathbf{R}_{\mathcal{T}_{r}}$ yields a scalar alignment score between each projection vector $\mathbf{p}$ and relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which then scales $\mathbf{R}_{\mathcal{T}_{r}}$ to generate an offset vector that modulates specific features in the entity embeddings and relation-meta. The updated embeddings ($\mathbf{h}_{j}^{\prime}$, $\mathbf{t}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$) thus capture task-specific nuances while maintaining consistency with the generalizable meta-knowledge. Although the three projection vectors could functionally be merged into a single perturbation term, maintaining separate projections for the head entity, relation-meta, and tail entity enhances modularity and flexibility. This design allows for more expressive or heterogeneous projection functions that can incorporate richer inductive biases or task-specific modeling needs in future extensions. Overall, this local adaptation scheme ensures that each task can adaptively focus on its unique aspects while preserving shared commonalities across related tasks.
To further optimize the adapted embeddings $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$, we define a scoring function in the projected space to measure the plausibility of each triplet $(h_{j},r,t_{j})$, which is given by:
$\mathrm{score}(h_{j},t_{j})=\left\lVert\mathbf{h}_{j}^{\prime}+\mathbf{R}_{\mathcal{T}_{r}}^{\prime}-\mathbf{t}_{j}^{\prime}\right\rVert_{2},$ (12)
where $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm. $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$ are the projected embeddings of the head entity, relation-meta, and tail entity, respectively. This formulation ensures that, in the task-specific projection space, the distance between the adapted head entity embedding $\mathbf{h}_{j}^{\prime}$ and the adapted tail entity embedding $\mathbf{t}_{j}^{\prime}$ is minimized after being translated by the relation-meta embedding $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$.
@@
$\mathcal{L}(\mathcal{S}_{r})=\sum_{(h_{j},r,t_{j})\in\mathcal{S}_{r}}\max\{0,\mathrm{score}(h_{j},t_{j})+\gamma-\mathrm{score}(h_{j},t^{\prime}_{j})\},$ (13)
where $\gamma$ is a margin hyperparameter, and $(h_{j},r,t^{\prime}_{j})$ represents a negative triplet generated by replacing the tail entity $t_{j}$ with a random entity $t^{\prime}_{j}$ such that $(h_{j},r,t^{\prime}_{j})\notin\mathcal{G}$.
@@
where $(h_{q},r,t^{\prime}_{q})$ is a negative triplet generated by corrupting the tail entity $t_{q}$ with a randomly sampled entity $t^{\prime}_{q}$. The optimization objective for training the entire model is to minimize $\mathcal{L}(\mathcal{Q}_{r})$.
@@ 5 Experiments
Experimental Setup. Following prior work, we set the embedding dimension to $100$ for Nell-One and FB15K-One, and $50$ for Wiki-One, initialized using TransE-pretrained weights provided by GMatching (Xiong et al., 2018). The number of neighbors per entity is capped at 50 for aggregation. For MoE, each expert is a two-layer MLP (hidden dimension: 64, ReLU). The gate network is also a two-layer MLP (hidden dimension: 64, output dimension: 1, and ReLU). We set the number of experts $M$ as $32$ and the number of selected experts $N$ as $5$. The margin $\gamma$ in Eq. 16 is set to $1$. MoEMeta is trained using Adam (batch size: $1,024$, learning rate: $0.001$). All results are averaged over 5 independent runs with different random seeds. All models are implemented in PyTorch and trained on a single Tesla P100 GPU. More reproducibility details are provided in Appendix B.1.

### v3

@@ 4.2 MoE-based Meta-Knowledge Learning
$\mathbf{r}_{i}$ $=\frac{1}{N}\sum_{j=1}^{M}g_{i,j}f_{j}(\mathbf{h}_{i},\mathbf{t}_{i};\bm{\theta}_{j}),$ (5)
$g_{i,j}$ $=\begin{cases}s_{i,j},&\text{if }s_{i,j}\in\text{Top}_{i}\bigl(\{s_{i,j}\mid 1\leq j\leq M\},N\bigr),\\ 0,&\text{otherwise},\end{cases}$ (6)
$s_{i,j}$ $=\text{softmax}(\text{Gate}(\mathbf{h}_{i},\mathbf{t}_{i};\bm{\theta}_{g})),$ (7)
where each $f_{j}(\cdot;\bm{\theta}_{j})$ is an expert network, and $M$ denotes the total number of experts. For each support triplet $i$, a gating network $\text{Gate}(\cdot;\bm{\theta}_{g})$ computes a relevance score $s_{i,j}$ for each expert $j$ conditioned on the corresponding input. These scores are normalized via a softmax function over all $M$ experts. The function $\text{Top}_{i}(\cdot,N)$ selects only the top $N$ highest scores among $M$ experts for each triplet $i$, setting the remaining scores to zero. This sparsity mechanism enables the model to selectively activate only a subset of experts for each support triplet, facilitating the learning of more specialized relational prototypes globally shared across tasks.
@@ 4.3 Task-Tailored Local Adaptation
Specifically, for each task, MoEMeta maintains a task-specific set of local parameters, denoted as $\bm{\eta}$, namely three randomly initialized projection vectors, $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$, which project the embeddings $\mathbf{h}_{i}$, $\mathbf{R}_{\mathcal{T}_{r}}$, and $\mathbf{t}_{i}$ into a task-specific subspace. Inspired by TransD (Ji et al., 2015), which addresses the limitations of fixed mapping strategies using per-entity and per-relation projections, our approach leverages task-specific projections to avoid overparametrization under few-shot settings. As all triplets within a task share the same relation, we adopt a unified projection vector $\mathbf{p}_{h}$ for all head entities, and $\mathbf{p}_{t}$ for all tail entities within the task. This provides a lightweight solution for enabling efficient task-specific adaptation with minimal parameter overhead.
@@
$\mathbf{h}_{i}^{\prime}$ $=\mathbf{h}_{i}+(\mathbf{p}_{h}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}},$ (9)
@@
$\mathbf{t}_{i}^{\prime}$ $=\mathbf{t}_{i}+(\mathbf{p}_{t}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}}.$ (11)
Above, the dot product $\mathbf{p}_{x}^{\top}\mathbf{R}_{\mathcal{T}_{r}}$ yields a scalar alignment score between each projection vector $\mathbf{p}_{x},x\in\{h,r,t\}$, and the relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$. This scalar score is then used to scale $\mathbf{R}_{\mathcal{T}_{r}}$, yielding an offset vector that modulates corresponding features in the entity embeddings and relation-meta. The updated embeddings ($\mathbf{h}_{i}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, $\mathbf{t}_{i}^{\prime}$) thus capture task-specific nuances while maintaining consistency with the generalizable meta-knowledge. Although the three projection vectors could functionally be merged into a single perturbation term, maintaining separate projections for the head entity, relation-meta, and tail entity enhances modularity and flexibility. This design allows for more expressive or heterogeneous projection functions that can incorporate richer inductive biases or task-specific modeling needs in future extensions. Overall, this local adaptation scheme ensures that each task can adaptively focus on its unique aspects while preserving shared commonalities across related tasks.
To further optimize the adapted embeddings $\mathbf{h}_{i}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{i}^{\prime}$, we define a scoring function in the projected space to measure the plausibility of each triplet $(h_{i},r,t_{i})$, which is given by:
$\mathrm{score}(h_{i},t_{i})=\left\lVert\mathbf{h}_{i}^{\prime}+\mathbf{R}_{\mathcal{T}_{r}}^{\prime}-\mathbf{t}_{i}^{\prime}\right\rVert_{2},$ (12)
where $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm. $\mathbf{h}_{i}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{i}^{\prime}$ are the projected embeddings of the head entity, relation-meta, and tail entity, respectively. This formulation ensures that, in the task-specific projection space, the distance between the adapted head entity embedding $\mathbf{h}_{i}^{\prime}$ and the adapted tail entity embedding $\mathbf{t}_{i}^{\prime}$ is minimized after being translated by the relation-meta embedding $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$.
@@
$\mathcal{L}(\mathcal{S}_{r})=\sum_{(h_{i},r,t_{i})\in\mathcal{S}_{r}}\max\{0,\mathrm{score}(h_{i},t_{i})+\gamma-\mathrm{score}(h_{i},t^{\prime}_{i})\},$ (13)
where $\gamma$ is a margin hyperparameter, and $(h_{i},r,t^{\prime}_{i})$ represents a negative triplet generated by replacing the tail entity $t_{i}$ with a randomly sampled entity $t^{\prime}_{i}\in\mathcal{E}$ such that $(h_{i},r,t^{\prime}_{i})\notin\mathcal{G}$.
@@
where $(h_{q},r,t^{\prime}_{q})$ is a negative triplet generated by corrupting the tail entity $t_{q}$ with a randomly sampled entity $t^{\prime}_{q}\in\mathcal{E}$. The optimization objective for training the entire model is to minimize $\mathcal{L}(\mathcal{Q}_{r})$.
@@ 5 Experiments
Experimental Setup. Following prior work, we set the embedding dimension to $100$ for Nell-One and FB15K-One, and $50$ for Wiki-One, initialized using TransE-pretrained weights provided by GMatching (Xiong et al., 2018). The number of neighbors per entity is capped at 50 for aggregation. For MoE, each expert network is a two-layer MLP (hidden dimension: 64, ReLU). The gating network is also a two-layer MLP (hidden dimension: 64, output dimension: 1, and ReLU). We set the number of experts $M$ as $32$ and the number of selected experts $N$ as $5$. The margin $\gamma$ in Eq. 16 is set to $1$. MoEMeta is trained using the Adam optimizer (batch size: $1,024$, learning rate: $0.001$). All results are averaged over 5 independent runs with different random seeds. All models are implemented in PyTorch and trained on a single Tesla P100 GPU. More details for reproducing our results are provided in Appendix B.1.